Based on the following information, what is the ROE?

Total assets: 100

Debt: 40

Sales: 80

Net income: 10

Select one:

a. 25%

b. 15.2%

c. 16.7%

d. 14.8%

e. 20%
Answer #1

Total assets=debt+equity

Hence equity=(100-40)=60

Hence ROE=Net income/equity

=(10/60)

which is equal to

=16.7%(Approx).